### v1

# Algebraic and symplectic curves of degree 8

Abstract.
We study the existence of some irreducible projective plane curves of degree $8$ with some prescribed topological type of singularities in the algebraic and symplectic worlds.
Partially supported by Grant PID2020-114750GB-C31 funded by MCIN/AEI/10.13039/ 501100011033 and Departamento de Ciencia, Universidad y Sociedad del Conocimiento of the Gobierno de Aragón (E22_20R: “Álgebra y Geometría”)
Introduction
Since the eighties the study of the theory of complex analytic and algebraic varieties have been enriched by the study of pseudo-holomorphic and complex symplectic varieties. In the case of curves in the projective plane these new objects are strongly related to braid monodromy, see [Orevkov2000, Kulikov:07, KhKu:03], and can be constructed by suitable deformations of some arrangements of algebraic plane curves.
The starting point of this paper is an unpublished idea of S.Yu. Orevkov, which is explained in [Golla-Starkston:22] and outlined in §2. The main idea is to deform symplectically a tricuspidal quartic (or deltoid) such that the tangent lines to the cusps are not concurrent. The idea of Orevkov, performed in detail by M. Golla and L. Starkston is to apply a standard Cremona transformation in order to obtain an irreducible symplectic curve with a configuration of topological type of singularities which does not exist in the algebraic category.
In this work, we replace the Cremona transformation by a Kummer cover, in order to compare the symplectic and algebraic structures of curves of degree $4n$ with $3n$ singularities having the topological type of $u^{2n}-v^{3}=0$. The case $n=2$ offers significant interesting properties and we focus our attention in this case since the study of algebraic structures seems to be cumbersome for $n>2$. We prove the existence of symplectic curves $C_{\text{\rm symp}}$ of degree $8$ with $6$ singular points of type $\mathbb{E}_{6}$.
For $n=2$ the primary goal is to determine all algebraic curves of degree $8$ with $6$ singular points with the topological of $\mathbb{E}_{6}$. Unfortunately, the goal was too ambitious and has not been reached. As it usually happens, the existence of symmetries is helpful and in this paper we determine all such curves fixed by a non-trivial projective automorphism. There is exactly one such curve $C_{8,2}$ (up to projective automorphism, of course) fixed by an involution and four such curves $C_{8,3}^{i}$, $i=1,\dots,r$, invariant by an automorphism of order $3$; there are no more curves invariants by automorphism. These four curves have equations in conjugate number fields $\mathbb{K}_{i}\subset\mathbb{C}$ isomorphic to $\mathbb{Q}[t]/p(t)$ where $p(t)$ is an irreducible polynomial of degree $4$. A main question is if they share topological properties. Two of the roots of $p(t)$ are real and two complex conjugate; in this last case, complex conjugation is a homeomorphism of $\mathbb{P}^{2}$ reversing orientations on the curves. In the general case, most likely these curves are rigid by dimension arguments.
Another result in this paper is that there is no homeomorphism of $\mathbb{P}^{2}$ sending $C_{\text{\rm symp}}$ to an algebraic symmetric curve, but it may be isotopic to a non symmetric one (if such a curve exists). There is also no homeomorphism of $\mathbb{P}^{2}$ sending $C_{8,2}$ to one of the $C_{8,3}^{i}$, and, besides complex conjugation, we do not know the existence of homeomorphism of $\mathbb{P}^{2}$ exchanging the curves $C_{8,3}^{i}$ (respecting or reversing orientations).
Some proofs need non-straightforward computer algebra steps and rely heavily on computations in Sagemath [Sagemath]. The steps are described in several notebooks located in https://github.com/enriqueartal/SymplecticOctics which can be executed either in a computer with the last version of Sagemath or online using Binder [binder].
In §1 we describe some known properties of the deltoid and compute a special presentation of the fundamental group of the complement of the deltoid and the tangent lines at the cusps. In §2 we study the topology of a symplectic deformation of the previous arrangement of curves.
This paper is inspirated in the ideas fruitfully discussed in the workshop Complex and symplectic curve configurations organized by M. Golla, P. Pokorav and L. Starkston in Nantes, December 2022, and in particular in [Golla-Starkston:22]. I would like to thank the organizers, the attendants, and specially the team work on braid monodromy and symplectic curves.
1. The deltoid and its tangents at the singular points
The deltoid (or tricuspidal quartic, i.e, plane quartic with three ordinary cusps) is an important plane projective curve. It is rigid, in the sense, that two deltoids are projectively isomorphic. As it is the dual of a nodal cubic, it has the following well-known property.
Property 1.1.
The three tangent lines at the cusps of a deltoid are concurrent lines.
A symmetric equation of the deltoid is
$y^{2}z^{2}+z^{2}x^{2}+x^{2}y^{2}-2xyz(x+y+z)=0.$
The equation of the curve in the right-hand side of Figure 1 is:
$v^{4}+4(1+u)v^{3}+18uv^{2}-27u^{2}=0.$
We are going to use this property to have a topological model of the curve and its tangents.
[…]
Figure 1. Left: usual deltoid. Right: real picture with real parts of non real branches.
Using the techniques of [ACCT], applied to the right-hand side of Figure 1, we obtain the following result. The justification of this figure can be found in the notebook ConstructionSymplecticGroup.
Proposition 1.2.
The braid monodromy of the deltoid projecting from the intersection point of the tangent line to the cusps, when one of these tangents is the line at infinity (as in Figure 1, right) is given by $(\sigma_{2}\cdot\sigma_{1})^{2}$ (for $L_{1}$) and $(\sigma_{2}\cdot\sigma_{3})^{2}$ (for $L_{2}$).
As a consequence, the fundamental group $G_{\Delta L}$ of the complement of the deltoid and the lines $L_{1},L_{2},L_{\infty}$ is generated by $c_{1},\dots,c_{4},\ell_{1},\ell_{2},\ell_{\infty}$ with the relations
  1. (R1)
    $[\ell_{2},c_{1}]=1$
  2. (R2)
    $\ell_{2}^{-1}\cdot c_{2}\cdot\ell_{2}=(c_{2}\cdot c_{3}\cdot c_{4})\cdot c_{3}\cdot(c_{2}\cdot c_{3}\cdot c_{4})^{-1}$
  3. (R3)
    $\ell_{2}^{-1}\cdot c_{3}\cdot\ell_{2}=(c_{2}\cdot c_{3})\cdot c_{4}\cdot(c_{2}\cdot c_{3})^{-1}$
  4. (R4)
    $\ell_{2}^{-1}\cdot c_{4}\cdot\ell_{2}=c_{2}$
  5. (R5)
    $\ell_{1}^{-1}\cdot c_{1}\cdot\ell_{1}=(c_{1}\cdot c_{2})\cdot c_{3}\cdot(c_{1}\cdot c_{2})^{-1}$
  6. (R6)
    $\ell_{1}^{-1}\cdot c_{2}\cdot\ell_{1}=(c_{1}\cdot c_{2})\cdot c_{1}\cdot(c_{1}\cdot c_{2})^{-1}$
  7. (R7)
    $\ell_{1}^{-1}\cdot c_{3}\cdot\ell_{1}=c_{1}\cdot c_{2}\cdot c_{1}^{-1}$
  8. (R8)
    $[\ell_{1},c_{4}]=1$
  9. (R9)
    $c\cdot\ell_{1}\cdot\ell_{2}\cdot\ell_{\infty}=1$
where $c=c_{1}\cdot\ldots\cdot c_{4}$.
This monodromy can be also computed using Sagemath [Sagemath] with the optional package Sirocco [Marco2016], but in this case it can be done directly.
In the presentation of $G_{\Delta L}$ we may omit the generator $\ell_{\infty}$ using (R9) which comes from the situation at infinity. Actually, Zariski-van Kampen method can be thought to happen in the blow-up of the projection point (the point at infinity of the vertical lines), see Figure 2. Then (R9) comes from the boundary of a neighbourhood of the exceptional divisor $E$, see [Mumford1961, ACMat:2020]. Note that the natural meridian $e$ of $E$ is the inverse of $c$. The normal crossing situation implies that $e$ (and hence $c$ and $\ell_{1}\cdot\ell_{2}\cdot\ell_{\infty}$) commute with $\ell_{1},\ell_{2},\ell_{\infty}$. This is a consequence of (R2)-(R7): $\ell_{1},\ell_{2}$ commute with $c$ as their conjugation action comes from braids.
[…]
Figure 2. Blow-up of the projection point.
Remark 1.3.
Note that $G_{\Delta L}$ is a semidirect product $\mathbb{F}_{4}\rtimes\mathbb{F}_{2}$ where $c_{1},\dots,c_{4}$ are the generators of normal subgroup $\mathbb{F}_{4}$, $\ell_{1},\ell_{2}$ are the generators of $\mathbb{F}_{4}$ and (R2)-(R7), determine the conjugation action.
This group has been computed in [ADT:10], but we need the above computation both for completeness and to deal with the symplectic deformations.
2. Symplectic deformations
In the context of symplectic geometry, it is possible to construct a deltoid for which the pseudo-holomorphic tangent lines at the cusps are not concurrent. This was communicated long time ago to the author by S.Yu. Orevkov and was formally written in [Golla-Starkston:22, § 8]. Moreover it can be done as a deformation of the algebraic curve which is an isotopy outside a neighbourhood of the triple point.
[…]
Figure 3. Symplectic deformation of an ordinary triple point.
Using the classical Seifert-van Kampen theorem, the fundamental group of the complement of this symplectic curve has the same presentation of the algebraic one, adding the relations from the situation in the right-hand side of Figure 3, i.e., $[\ell_{1},\ell_{2}]=[\ell_{1},\ell_{\infty}]=[\ell_{2},\ell_{\infty}]=1$. This technique has been used in [ACM:2020b, AC:1998, CatWaj:05] Actually, the following holds.
Corollary 2.1.
The fundamental group $G_{s\Delta L}$ of the complement of the symplectic deltoid and the tangent lines at the cusps has the generators and relators of Proposition 1.2 plus the relation
  1. (R10)
    $[\ell_{1},\ell_{2}]=1$.
It is useful to have a semidirect presentation of this group.
Corollary 2.2.
The group $G_{s\Delta L}$ is a semidirect product $G_{0}\rtimes\mathbb{Z}^{2}$ where the action is as in the algebraic case and
$G_{0}=\langle c_{1},\dots,c_{4}\mid c_{3}\cdot c_{4}\cdot c_{3}=c_{4}\cdot c_{3}\cdot c_{4},\ c_{1}\cdot c_{2}\cdot c_{1}=c_{2}\cdot c_{1}\cdot c_{2}\ \rangle.$
Proof.
We start with the semidirect product structure $G_{\Delta L}=\mathbb{F}_{4}\rtimes\mathbb{F}_{2}$ and the natural epimorphism $G_{\Delta L}\twoheadrightarrow G_{s\Delta L}$: