### v2

# Algebraic and symplectic curves of degree 8

Abstract.
We study the existence of some irreducible projective plane curves of degree $8$ with some prescribed topological type of singularities in the algebraic and symplectic worlds.
Partially supported by Grant PID2020-114750GB-C31 funded by MCIN/AEI/10.13039/ 501100011033 and Departamento de Ciencia, Universidad y Sociedad del Conocimiento of the Gobierno de Aragón (E22_20R: “Álgebra y Geometría”)
Introduction
Since the eighties the study of the theory of complex analytic and algebraic varieties have been enriched by the study of pseudo-holomorphic and complex symplectic varieties. In the case of curves in the projective plane these new objects are strongly related to braid monodromy, see [Ore00, Kul07, KK03], and can be constructed by local deformations of arrangements of algebraic plane curves which can be expressed in terms of braid monodromy factorizations which are locally algebraic.
The starting point of this paper is an unpublished idea of S.Yu. Orevkov, which is explained in [GS22] and outlined in §2. The main idea is to deform symplectically a tricuspidal quartic (or deltoid) such that the tangent lines to the cusps are not concurrent. The idea of Orevkov, performed in detail by M. Golla and L. Starkston is to apply a standard Cremona transformation in order to obtain an irreducible symplectic curve with a configuration of topological type of singularities which does not exist in the algebraic category.
In this work, we replace the Cremona transformation by a Kummer cover, in order to compare the symplectic and algebraic structures of curves of degree $4n$ with $3n$ singularities having the topological type of $u^{2n}-v^{3}=0$. The case $n=2$ offers significant interesting properties and we focus our attention in this case since the study of algebraic structures seems to be cumbersome for $n>2$. We prove the existence of symplectic curves $C_{\text{\rm symp}}$ of degree $8$ with $6$ singular points of type $\mathbb{E}_{6}$.
For $n=2$ the primary goal is to determine all algebraic curves of degree $8$ with $6$ singular points with the topological type of $\mathbb{E}_{6}$. Unfortunately, the goal was too ambitious and has not been reached. As it usually happens, the existence of symmetries is helpful and in this paper we determine all such curves fixed by a non-trivial projective automorphism. There is exactly one such curve $C_{8,2}$ (up to projective automorphism, of course) fixed by an involution and four such curves $C_{8,3}^{i}$, $i=1,\dots,r$, invariant by an automorphism of order $3$; there are no more curves invariant by automorphism. These four curves have equations in conjugate number fields $\mathbb{K}_{i}\subset\mathbb{C}$ isomorphic to $\mathbb{Q}[t]/p(t)$ where $p(t)$ is an irreducible polynomial of degree $4$. A main question is if they share topological properties. Two of the roots of $p(t)$ are real and two complex conjugate; in this last case, complex conjugation is a homeomorphism of $\mathbb{P}^{2}$ reversing orientations on the curves. In the general case, most likely these curves are rigid by dimension arguments.
Another result in this paper is that there is no homeomorphism of $\mathbb{P}^{2}$ sending $C_{\text{\rm symp}}$ to an algebraic symmetric curve, but it may be isotopic to a non symmetric one (if such a curve exists). There is also no homeomorphism of $\mathbb{P}^{2}$ sending $C_{8,2}$ to one of the $C_{8,3}^{i}$, and, besides complex conjugation, we do not know the existence of homeomorphism of $\mathbb{P}^{2}$ exchanging the curves $C_{8,3}^{i}$ (respecting or reversing orientations).
Some proofs need non-straightforward computer algebra steps and rely heavily on computations in Sagemath [Ste23]. The steps are described in several notebooks located in https://github.com/enriqueartal/SymplecticOctics which can be executed either in a computer with the last version of Sagemath or online using Binder [PJe18].
In §1 we describe some known properties of the deltoid and compute a special presentation of the fundamental group of the complement of the deltoid and the tangent lines at the cusps. In §2 we study the topology of a symplectic deformation of the previous arrangement of curves.
This paper is inspirated in the ideas fruitfully discussed in the workshop Complex and symplectic curve configurations organized by M. Golla, P. Pokorav and L. Starkston in Nantes, December 2022, and in particular in [GS22]. I would like to thank the organizers, the attendants, and specially the team work on braid monodromy and symplectic curves.
1. The deltoid and its tangents at the singular points
The deltoid (or tricuspidal quartic, i.e, plane quartic with three ordinary cusps) is an important plane projective curve. It is rigid, in the sense, that two deltoids are projectively isomorphic. As it is the dual of a nodal cubic, it has the following well-known property.
Property 1.1.
The three tangent lines at the cusps of a deltoid are concurrent lines.
A symmetric equation of the deltoid is
$y^{2}z^{2}+z^{2}x^{2}+x^{2}y^{2}-2xyz(x+y+z)=0.$
The equation of the curve in the right-hand side of Figure 1 is
$v^{4}+4(1+u)v^{3}+18uv^{2}-27u^{2}=0;$
the line at infinity is the tangent line to one of the cusps; the other cusps are $(0,0),(1,-3)$ and they have vertical tangent lines. The vertical lines $u=a$, $a\in\mathbb{R}$, intersect the real part of the curve at two real points (solid curves in the right-hand side of Figure 1) and at two other points $(a,v_{0}(a)\pm\sqrt{-1}v_{1}(a))$, $v_{0}(a)\in\mathbb{R}$, $v_{1}(a)\in\mathbb{R}_{>0}$; the dotted curve in the right-hand side of Figure 1 represents $(a,v_{0}(a))$. This picture provides a topological model of the curve and its tangents.
[…]
Figure 1. Left: usual deltoid. Right: real picture with real parts of non real branches.
Using the techniques of [ACCT01], applied to the right-hand side of Figure 1, we obtain the following result. The justification of this figure can be found in the notebook ConstructionSymplecticGroup.
$c_{1}$$c_{2}$$c_{3}$$c_{4}$$c$$(u_{0},v_{0})$$u=u_{0}$$\ell_{1}$$\ell_{2}$$\ell_{\infty}$$(u_{0},v_{0})$$v=v_{0}$
Figure 2. Generators of the presentation in Proposition 1.2. The base point is $(u_{0},v_{0})$, where $10<u_{0}<$ 1 (the $u$-coordinates) of the affine singular points, and $v_{0}\gg 0$.
Proposition 1.2.
The braid monodromy of the deltoid projecting from the intersection point of the tangent line to the cusps, when one of these tangents is the line at infinity (as in Figure 1, right) is given by $(\sigma_{2}\cdot\sigma_{1})^{2}$ (for $L_{1}$) and $(\sigma_{2}\cdot\sigma_{3})^{2}$ (for $L_{2}$).
As a consequence, the fundamental group $G_{\Delta L}$ of the complement of the deltoid and the lines $L_{1},L_{2},L_{\infty}$ is generated by $c_{1},\dots,c_{4},\ell_{1},\ell_{2},\ell_{\infty}$ with the relations
  1. (R1)
    $[\ell_{2},c_{1}]=1$
  2. (R2)
    $\ell_{2}^{-1}\cdot c_{2}\cdot\ell_{2}=(c_{2}\cdot c_{3}\cdot c_{4})\cdot c_{3}\cdot(c_{2}\cdot c_{3}\cdot c_{4})^{-1}$
  3. (R3)
    $\ell_{2}^{-1}\cdot c_{3}\cdot\ell_{2}=(c_{2}\cdot c_{3})\cdot c_{4}\cdot(c_{2}\cdot c_{3})^{-1}$
  4. (R4)
    $\ell_{2}^{-1}\cdot c_{4}\cdot\ell_{2}=c_{2}$
  5. (R5)
    $\ell_{1}^{-1}\cdot c_{1}\cdot\ell_{1}=(c_{1}\cdot c_{2})\cdot c_{3}\cdot(c_{1}\cdot c_{2})^{-1}$
  6. (R6)
    $\ell_{1}^{-1}\cdot c_{2}\cdot\ell_{1}=(c_{1}\cdot c_{2})\cdot c_{1}\cdot(c_{1}\cdot c_{2})^{-1}$
  7. (R7)
    $\ell_{1}^{-1}\cdot c_{3}\cdot\ell_{1}=c_{1}\cdot c_{2}\cdot c_{1}^{-1}$
  8. (R8)
    $[\ell_{1},c_{4}]=1$
  9. (R9)
    $c\cdot\ell_{1}\cdot\ell_{2}\cdot\ell_{\infty}=1$
where $c=c_{1}\cdot\ldots\cdot c_{4}$.
This monodromy can be also computed using Sagemath [Ste23] with the optional package Sirocco [MR16], but in this case it can be done directly.
In the presentation of $G_{\Delta L}$ we may omit the generator $\ell_{\infty}$ using (R9) which comes from the situation at infinity. Actually, Zariski-van Kampen method can be thought to happen in the blow-up of the projection point (the point at infinity of the vertical lines), see Figure 3. Then (R9) comes from the boundary of a neighbourhood of the exceptional divisor $E$, see [Mum61, ACM20c]. Note that the natural meridian $e$ of $E$ is the inverse of $c$. The normal crossing situation implies that $e$ (and hence $c$ and $\ell_{1}\cdot\ell_{2}\cdot\ell_{\infty}$) commute with $\ell_{1},\ell_{2},\ell_{\infty}$. This is a consequence of (R2)-(R7): $\ell_{1},\ell_{2}$ commute with $c$ as their conjugation action comes from braids.
[…]
Figure 3. Blow-up of the projection point.
Remark 1.3.
Note that $G_{\Delta L}$ is a semidirect product $\mathbb{F}_{4}\rtimes\mathbb{F}_{2}$ where $c_{1},\dots,c_{4}$ are the generators of normal subgroup $\mathbb{F}_{4}$, $\ell_{1},\ell_{2}$ are the generators of $\mathbb{F}_{2}$ and (R2)-(R7), determine the conjugation action.
This group has been computed in [ADT10], but we need the above computation both for completeness and to deal with the symplectic deformations.
2. Symplectic deformations
In the context of symplectic geometry, it is possible to construct a deltoid for which the pseudo-holomorphic tangent lines at the cusps are not concurrent. This was communicated long time ago to the author by S.Yu. Orevkov and was formally written in [GS22, § 8]. Moreover it can be done as a deformation of the algebraic curve which is an isotopy outside a neighbourhood of the triple point.
[…]
Figure 4. Symplectic deformation of an ordinary triple point.
Using the classical Seifert-van Kampen theorem, the fundamental group of the complement of this symplectic curve has the same presentation of the algebraic one, adding the relations from the situation in the right-hand side of Figure 4, i.e., $[\ell_{1},\ell_{2}]=[\ell_{1},\ell_{\infty}]=[\ell_{2},\ell_{\infty}]=1$. This technique has been used in [ACM20b, AC98, CW05] Actually, the following holds.
Corollary 2.1.
The fundamental group $G_{s\Delta L}$ of the complement of the symplectic deltoid and the tangent lines at the cusps has the generators and relators of Proposition 1.2 plus the relation
  1. (R10)
    $[\ell_{1},\ell_{2}]=1$.
It is useful to have a semidirect presentation of this group.
Corollary 2.2.
The group $G_{s\Delta L}$ is a semidirect product $G_{0}\rtimes\mathbb{Z}^{2}$ where the action is as in the algebraic case and
$G_{0}=\langle c_{1},\dots,c_{4}\mid c_{3}\cdot c_{4}\cdot c_{3}=c_{4}\cdot c_{3}\cdot c_{4},\ c_{1}\cdot c_{2}\cdot c_{1}=c_{2}\cdot c_{1}\cdot c_{2}\ \rangle.$
Proof.
We start with the semidirect product structure $G_{\Delta L}=\mathbb{F}_{4}\rtimes\mathbb{F}_{2}$ and the natural epimorphism $G_{\Delta L}\twoheadrightarrow G_{s\Delta L}$:
[…]
The semidirect structure comes from the fact that the below exact sequence splits as seen in the relations. The conjugation in the first exact sequence is given by the action of the braids $\tau_{1}:=(\sigma_{2}\cdot\sigma_{1})^{2}$ (for $\ell_{1}$) and $\tau_{2}:=(\sigma_{2}\cdot\sigma_{3})^{2}$ (for $\ell_{2}$). Since $\ell_{1},\ell_{2}$ commute in $G_{s\Delta L}$, then in $G_{0}$ we have the relations (checked in ConstructionSymplecticGroup)
$c_{i}^{\tau_{1}\cdot\tau_{2}}=c_{i}^{\tau_{2}\cdot\tau_{1}},\qquad i=1,\dots,4,$
which translates into the relations in the statement. ∎
3. Algebraic and symplectic Cremona transformations
Following the ideas of Orevkov, Golla and Starkston formalized in [GS22, § 8] an example of rational singular curves which exist in the symplectic category and not in the algebraic one.
The most well known birational is the map
${\mathbb{P}^{2}}$${\mathbb{P}^{2}}$${{[x:y:z]}}$${{[yz:zx:xy]}.}$
Geometrically is obtained by the blow-up of the points [1:0:0],[0:1:0],[0:0:1][1:0:0],[0:1:0],[0:0:1][ 1 : 0 : 0 ] , [ 0 : 1 : 0 ] , [ 0 : 0 : 1 ] and the blow-down of the strict transforms of the lines $x=0,y=0,z=0$ which are pairwise disjoint smooth rational $(-1)$-curves. We can also consider a symplectic Cremona transformation which gives the following result.
Proposition 3.1 ([GS22, § 8]).
Let $\Sigma_{\text{\rm alg}}$ (resp. $\Sigma_{\text{\rm symp}}$) be the space of algebraic (resp. symplectic) irreducible curves of degree $8$ in $\mathbb{P}^{2}$ having three singular points with the topological type of $u(v^{3}+u^{5})=0$.
  1. (1)
    The space $\Sigma_{\text{\rm alg}}$ is empty.
  2. (2)
    The space $\Sigma_{\text{\rm symp}}$ is non-empty and it can be embedded in the space of symplectic deltoids such that their tangent lines to the cuspidal points are not concurrent.
We can go further and compute some topological invariants of this curve, in particular the fundamental group of its complement.
Corollary 3.2.
If $C\in\Sigma_{\text{\rm symp}}$ comes from a Cremona transformation associated to the tangent lines of a symplectic deltoid (isotopic to an algebraic one), then its fundamental group is the non-abelian semidirect product $\mathbb{Z}/3\rtimes\mathbb{Z}/8$.
Proof.
If $P$ is an ordinary double point and two commuting meridians of the branches, a meridian of the exceptional component of the blow-up of the $P$ is the product of the meridians, see e.g. [ACM20a, Lemma 3.6] (probably well-known result).
The complement of $C$ is homeomorphic to the complement of the strict transform of the deltoid and the tangent lines by the blow-ups. For the total transform, we have to add the exceptional components. From the deformation in Figure 4, we see that these meridians are $\ell_{1}\cdot\ell_{2}$, $\ell_{1}\cdot\ell_{\infty}$, and $\ell_{2}\cdot\ell_{\infty}$.
From [Fuj82, Lemma 4.18], the fundamental group of the complement of the strict transform is obtained by killing these meridians. These new relations are summarized in
$\ell:=\ell_{1}=\ell_{2}=\ell_{\infty},\quad\ell^{2}=1$
and clearly imply (R10). The relation (R9) becomes $c=\ell$. Relations (R1) and (R8) become $[\ell,c_{1}]=[\ell,c_{2}]$; since (R4) becomes $c_{2}=c_{4}$ we obtain also that $\ell$ is central. From (R5) we can eliminate $c_{1}$ and from a simple computation we obtain that $c_{2},c_{3}$ generate with the relations
$c_{2}\cdot c_{3}\cdot c_{2}=c_{3}\cdot c_{2}\cdot c_{3},\quad c_{2}^{2}\cdot c_{3}^{2}\text{ central and of order }2.$
The normal subgroup of order $3$ is generated by $c_{2}\cdot c_{3}^{-1}=(c_{2}\cdot c_{3})^{4}$ and the subgroup of order $8$ is generated by $c_{2}$. ∎
Remark 3.3.
This group is also the fundamental group of the complement of an algebraic curve, as it is shown using similar techniques in [Ulu01].
4. Kummer covers
With the same ideas of §3, we are going to construct new examples replacing the standard Cremona transformation by Kummer covers, i.e., Galois covers
${\mathbb{P}^{2}}$${\mathbb{P}^{2}}$${{[x:y:z]}}$${{[x^{n}:y^{n}:z^{n}]}.}$
Starting from three pseudo-holomorphic non-concurrent lines there is a symplectic counterpart.
Let us recall that an $\mathbb{E}_{6}$-singularity is a germ of plane curve singularity in a smooth surface isomorphic to $u^{3}-v^{4}=0$ in $(\mathbb{C}^{2},0)$ (with local coordinates $u,v$).
Proposition 4.1.
There are irreducible symplectic curves $C_{\text{\rm symp}}$ of degree $8$ in $\mathbb{P}^{2}$ with $6$ singular points of type $\mathbb{E}_{6}$ for which the fundamental group $G_{\text{\rm symp}}$ of their complement is generated by $c^{\prime}_{1},c_{2},c_{3},c_{4}$, $c^{\prime}_{1}=c_{2}^{-1}\cdot c_{1}\cdot c_{2}$, with relations
$[c_{2},c_{4}]=[c^{\prime}_{1},c_{3}]=1,\ c^{\prime}_{1}\cdot c_{2}\cdot c^{\prime}_{1}=c_{2}\cdot c^{\prime}_{1}\cdot c_{2},\ c_{3}\cdot c_{2}\cdot c_{3}=c_{2}\cdot c_{3}\cdot c_{2},$
$c_{3}\cdot c_{4}\cdot c_{3}=c_{4}\cdot c_{3}\cdot c_{4},\ (c_{2}\cdot c^{\prime}_{1}\cdot c_{3}\cdot c_{4})^{2}=1,$
and the conjugation action is derived from the action in Remark 1.3.
As in §3, we denote by $\Lambda_{\text{\rm symp}}$ and $\Lambda_{\text{\rm alg}}$ the spaces of symplectic or algebraic curves of degree $8$ having $6$ singular points of type $\mathbb{E}_{6}$.
Proof.
The existence of such a curve $C_{\text{\rm symp}}$ comes from a symplectic Kummer cover for $n=2$, starting from a symplectic deltoid as in §2, taking the tangent lines to the cusps for the ramification lines of the Kummer cover. The degree of the preimage of the deltoid is $8$ and each cusp produces two $\mathbb{E}_{6}$ points.
For the fundamental group, let $G_{\textrm{orb22}}$ be the orbifold fundamental group of the complement of the deltoid where the orbifold structure comes from the action of $\mathbb{Z}/2\times\mathbb{Z}/2$ as deck group of the Kummer cover.
Hence, $G$ is the quotient of the group in Corollary 2.1 with some extra relations
  1. (R11)
    $\ell_{1}^{2}=1$
  2. (R12)
    $\ell_{2}^{2}=1$
  3. (R13)
    $\ell_{\infty}^{2}=1$
As in the proof of Corollary 2.2, this group $G$ is a semidirect product $G_{\text{\rm symp}}\rtimes\mathbb{Z}/2\times\mathbb{Z}/2$. In order to find $G_{\text{\rm symp}}$ we consider the relations $c_{i}^{\tau_{j}}=c_{i}^{\tau_{j}^{-1}}$, for $i=1,\dots,4$ and $j=1,2$. Moreover we can combine the relations (R9) and (R13) to rewrite them in terms only of $c_{1},\dots,c_{4}$. Replacing $c_{1}$ by $c_{1}^{\prime}$ we obtain the relation of the statement. Details can be found in ConstructionSymplecticGroup.
Since the fundamental group of the complement of $C_{\text{\rm symp}}$ is the kernel of the epimorphism $G\to\mathbb{Z}/2\times\mathbb{Z}/2$ given by
$c_{i}\mapsto(0,0),\quad\ell_{1}\mapsto(1,0),\quad\ell_{2}\mapsto(0,1),\quad\ell_{\infty}\mapsto(1,1),$
we obtain that this group is $G_{\text{\rm symp}}$. ∎
Remark 4.2.
Using GAP4 [GAP22] via Sagemath [Ste23] we have:
$G/G^{\prime}\cong\mathbb{Z}/8,\quad G^{\prime}/G^{\prime\prime}\cong\mathbb{Z}/3,\quad G^{\prime\prime}/G^{\prime\prime\prime}\cong(\mathbb{Z}/2)^{6},\quad G^{\prime\prime\prime}/G^{(4)}\cong\mathbb{Z}^{9}\oplus(\mathbb{Z}/2)^{5}\oplus\mathbb{Z}/4.$
We need to understand $\Lambda_{\text{\rm alg}}$ in order to check if the elements found in $\Lambda_{\text{\rm symp}}$ are isotopic to algebraic curves. Unfortunately computations are cumbersome and our attempts failed. Most probably this space is discrete, and we have been able to obtain some particular elements. Some geometric properties of these curves are presented in the following section.
5. Symmetries and other properties of curves in $\Lambda_{\text{\rm alg}}$
We want to study the properties of the curves in $\Lambda_{\text{\rm alg}}$. We start with the symmetry properties. Let us recall that the automorphism group of $\mathbb{P}^{2}$ is the group $\operatorname{PGL}(3;\mathbb{C})$. The elements of finite order correspond to diagonalizable matrices (up to scalar multiplication) whose eigenvalues are roots of unity.
Example 5.1.
The involutions of $\mathbb{P}^{2}$ correspond to matrices which are conjugate to the diagonal matrix $(1,1,-1)$, i.e. conjugate to the automorphism $\Phi:\mathbb{P}^{2}\to\mathbb{P}^{2}$ such that $\Phi([x:y;z])=[x:y:-z]$. These automorphisms have an isolated fixed point, $[0:0:1]$ (for the eigenspace of dimension $1$), and a line of fixed points, $z=0$ (for the eigenspace of dimension $2$).
The quotient of $\mathbb{P}^{2}$ by an involution is isomorphic to the weighted projective plane ($\mathbb{P}^{2}_{(1,1,2)}$. Let $\omega:=(p,q,r)$ be a positive integer vector with pairwise coprime weights and consider the weighted projective plane $\mathbb{P}^{2}_{\omega}$, see [Dol82] for details. It is a normal projective surface structure in the quotient
$\mathbb{C}^{3}\setminus\{\mathbf{0}\}/(x,y,z)\sim(t^{p}x,t^{q}y,t^{r}z);$
its elements are denoted $[x:y:z]_{\omega}$. The curves are the zero loci of $\omega$-weighted homogeneous polynomials and since they are in general Weil divisors. Bézout’s formula is also valid in the weighted projective planes. Namely if $C_{1},C_{2}$ are curves defined by $\omega$-weighted homogeneous polynomials of degrees $d_{1},d_{2}$ respectively. Then
(5.1) $C_{1}\cdot C_{2}=\frac{d_{1}\cdot c_{2}}{p\cdot q\cdot r}.$
Lemma 5.2.
Let $C\in\Lambda_{\text{\rm alg}}$ symmetric by the action of a projective involution $\Phi_{2}$. Then two of the singular points are in the line of fixed points in $\Phi_{2}$ and the other ones form two orbits.
Proof.
We can assume that $\Phi_{2}([x:y:z])=[x:y:-z]$; let $F_{8}(x,y,z)=0$ be the equation of $C$. Since the curve is invariant, we have that $F_{8}(x,y,z)=F_{8}(x,y,-z)$, i.e., $F_{8}(x,y,z)=G_{8}(x,y,z^{2})$ where $G_{8}$ is a $(1,1,2)$-weighted homogeneous polynomial of degree $8$. The quotient $\tilde{C}$ of $C$ is a curve in ($\mathbb{P}^{2}_{(1,1,2)}$ with equation $G_{8}(x_{2},y_{2},z_{2})=0$.
An $\mathbb{E}_{6}$ point cannot be the isolated fixed point $[0:0:1]$ of $\Phi_{2}$. Let us assume that no singular point is in the line of fixed points. Then, the quotient of $C$ in ($\mathbb{P}^{2}_{(1,1,2)}$ is a curve of degree $8$ with three triple points of type $\mathbb{E}_{6}$. There is no line $L_{1}$ of equation $ax_{2}+by_{2}=0$ through two singular points. If it would be the case, since $L_{1}$ is of degree $1$, we would have
$4=\frac{\deg L_{1}\cdot\deg\tilde{C}}{2}=L_{1}\cdot\tilde{C}\geq 3+3,$
so it is not possible. It is not difficult to check that three points in ($\mathbb{P}^{2}_{(1,1,2)}$ such that no pair is contained in a line, are contained in a curve $C_{2}$ of degree $2$. Then:
$8=\frac{\deg C_{2}\cdot\deg\tilde{C}}{2}=C_{2}\cdot\tilde{C}\geq 3+3+3,$
which is also impossible. The only possible case is the one in the statement. ∎
Example 5.3.
There are two types of automorphisms $\mathbb{P}^{2}$ of order $3$. The first one corresponds to matrices which are conjugate to the diagonal matrix $(1,1,-\zeta)$, where $\zeta:=\exp\frac{2\sqrt{-1}\pi}{3}$, with one isolated fixed point and a line of fixed points.
The second type corresponds to matrices which are conjugate to the diagonal matrix $(\zeta,\overline{\zeta},1)$ and has three isolated fixed points. There are exactly three fixed lines, the lines joining the fixed points.
Lemma 5.4.
Let $C\in\Lambda_{\text{\rm alg}}$ symmetric by the action of a projective automorphism $\Phi_{3}$ of order $3$. Then $\Phi_{3}$ has no line of fixed points, there are $2$ orbits and the curve passes through two isolated fixed points of $\Phi_{3}$ (tangent to the fixed lines not containing the two fixed points in the curve).
Proof.
Let us suppose first that $\Phi_{3}$ has a line of fixed points. At most two singular points can be in this line by Bézout Theorem, but actually none of them can be in the line since the orbits have one or three elements. But the points in the orbits are aligned which is contradiction again with Bézout Theorem.
Hence $\Phi_{3}$ has three fixed points, say $P_{1},P_{2},P_{3}$. These points cannot be singular points of the curve since an $\mathbb{E}_{6}$ cannot be fixed an isolated fixed point of an action of order $3$.
Hence, the singular points form two orbits. Let us consider the lines joining the fixed points, say $L_{i}$ is the line joining $P_{j}$ and $P_{k}$, $\{i,j,k\}=\{1,2,3\}$. Since the action is free on $L_{i}\setminus\{P_{j},P_{k}\}$, it must intersect $C$ with intersection number $6$. This is only achieved (after reordering) if $P_{1},P_{2}\in C$, $P_{3}\notin C$, $L_{2}$ is tangent to $C$ at $P_{1}$ and $L_{1}$ is tangent to $C$ at $P_{2}$. ∎
Example 5.5.
There are several types of automorphisms of order $n>3$, depending on the different configurations of eigenvalues.
Lemma 5.6.
There is no $C\in\Lambda_{\text{\rm alg}}$ symmetric by the action of a projective automorphism $\Phi$ of order $n>3$.
Proof.
Note first that we cannot have a line of fixed points, only isolated points. The case $n>7$ is ruled out immediately.
For $n=4$, there are two possible types of automorphisms $\Phi_{4}$, conjugate to the diagonal matrices of either $(\sqrt{-1},1,1)$ or $(\sqrt{-1},-1,1)$. The first case (with one isolated fixed point and a line of fixed points) is ruled as in the first part of the proof of Lemma 5.4. For the second case, we have three isolated fixed points of order $4$. The line joining two of them, say $P_{1},P_{2}$, is a line of fixed points for $\Phi_{4}^{2}$. The points $P_{i}$ cannot be singular points of the curve $C$. The only possible option is to have an orbit of four singular points and another one of two points. But the orbit of four points is formed by aligned points and it is forbidden by Bézout Theorem.
In the case $n=5$, let $\Phi_{5}$ be such a automorphism. If there is a line of fixed points we conclude again as in the first part of the proof of Lemma 5.4. Let us assume that there are three isolated fixed points, which cannot be singular in the curve. The set singular points must be the union of orbits of $5$ elements, which is not possible.
For the case $n=6$, the restrictions for $n=2,3$ give only one possible case, corresponding to an automorphism $\Phi_{6}$ conjugate to a diagonal matrix $(-\zeta,-\overline{\zeta},1)$, hence only three fixed points which cannot be singular points. The singular points form one orbit; then for $\Phi_{6}^{3}$ we would have three orbits of two points which has been ruled out in Lemma 5.2. ∎
These curves have interesting properties from the birational point of view. Let $C\in\Lambda_{\text{\rm alg}}$ (though most of the following facts may be also valid in the symplectic case). Le $P_{1},\dots,P_{6}$ be the singular points. They are not in a conic (from Bézout Theorem). Let $\mathcal{C}_{i}$, $1\leq i\leq 6$, be the unique conic passing through $P_{1},\dots,\widehat{P_{i}},\dots,P_{6}$. Again, by Bézout Theorem, these conics are irreducible.
Proposition 5.7.
Let $\Psi:\mathbb{P}^{2}\dashrightarrow\mathbb{P}^{2}$ the birational map obtained by blowing-up the points $P_{1},\dots,P_{6}$ and blowing-down the strict transforms of $\mathcal{C}_{1},\dots,\mathcal{C}_{6}$. Let $Q_{1},\dots,Q_{6}$ be the images of the conics and let $\mathcal{D}_{1},\dots,\mathcal{D}_{6}$ the images of the exceptional components.
Then, the strict transform of $C$ is a smooth quartic curve $D$ passing through the points $Q_{1},\dots,Q_{6}$. There exist six points $R_{1},\dots,R_{6}$ such that as divisors
$D\cdot\mathcal{D}_{i}=Q_{1}+\dots+\widehat{Q_{i}}+\dots+Q_{6}+3R_{i}.$
These twelve points are pairwise distinct.
Note that in particular, $C$ is not hyperelliptic.
Proof.
The map $\Psi$ factors as in the following diagram
${X}$${\mathbb{P}^{2}}$${\mathbb{P}^{2}}$$\scriptstyle{\sigma_{1}}$$\scriptstyle{\sigma_{2}}$$\scriptstyle{\Psi}$
The map $\sigma_{1}$ is the composition of the blow-ups of the points $P_{1},\dots,P_{6}$. Under these blow-ups, let us denote by $\mathcal{D}_{i}$ the exceptional divisors, and denote also by $\mathcal{C}_{i}$ the strict transform of $\mathcal{C}_{i}$. As each conic has been affected by $5$ blow-ups, $(\mathcal{C}_{i})^{2}_{X}=-1$, and these strict transforms are pairwise disjoint. Hence the map $\sigma_{2}$ is the blow-down of the curves $\mathcal{C}_{i}$. Under these blow-downs, the images $\mathcal{D}_{i}=\sigma_{2}(\mathcal{D}_{i})$ are conics; they pass through $5$ of the six exceptional points $Q_{j}=\sigma_{2}(\mathcal{C}_{j})$.
The other interesection point is the strict transform of a singular point which becomes a smooth point after blowing-up having intersection number $3$ with the exceptional divisor. ∎
Unfortunately, this description is not useful for the computations.
6. Algebraic curves with $\mathbb{Z}/2$-action
From the lemmas in §5, we can assume that $C_{8,2}\in\Lambda_{\text{\rm alg}}$ is fixed by the involution $\Phi_{2}:\mathbb{P}^{2}\to\mathbb{P}^{2}$ given by $\Phi_{2}([x:y:z])=[x:y:-z]$ and that two of the $\mathbb{E}_{6}$ points are $P_{1}=[1:0:0]$ and $P_{2}=[0:1:0]$. The isolated fixed point $[0:0:1]$ is not in the curve. The tangent lines to $P_{i}$ must be fixed by the action and it is easily seen that they are not tangent to $z=0$, hence the tangent lines are $L_{x}:\{x=0\}$ and $L_{y}:\{y=0\}$.
The quotient $\mathbb{P}^{2}/\Phi_{2}$ is isomorphic to the weighted projective plane $\mathbb{P}^{2}_{\omega}$, $\omega=(1,1,2)$, and the map is $\pi:\mathbb{P}^{2}\to\mathbb{P}^{2}_{\omega}$ where $\pi([x:y:z])=[x:y:z^{2}]_{\omega}$. From the orbifold point of view there is an orbifold $X_{2}$ constructed on $\mathbb{P}^{2}_{\omega}$ with the usual orbifold structure around $[0:0:1]_{\omega}$ and also on the line $L_{z}:\{z=0\}$, with an action of the cyclic group of order $2$.
Lemma 6.1.
Let $\tilde{C}_{8,2}:=\Phi_{2}(C_{8,2})$. Then $\tilde{C}_{8,2}$ is a curve of $\omega$-degree $8$, with two singular points $\mathbb{E}_{6}$ and two ordinary cusps (not two of them in the same curve of $\omega$-degree $1$). Moreover there is a curve of $\omega$-degree $2$ tangent to the two cusps.
This is obvious from the description of $C_{8,2}$. Note that the two cusps come from singular points in the line of fixed points, so they are not in a curve of $\omega$-degree $1$; for any other pair of points, the fact that two singular points are not on a curve of $\omega$-degree $1$ follows immediately from (5.1).
The way to compute the space of all such curves (up to automorphism) is the following one. We start with a polynomial
$f(x,y,z)=\sum_{i+j+2k=8}a_{ijk}x^{i}y^{j}z^{k}.$
Since it does not pass through $[0:0:1]_{\omega}$, we may assume that $a_{004}=1$. Recall that
$\operatorname{Aut}\mathbb{P}^{2}_{\omega}=\left\{\Phi_{B,c}\mid\right.\left|B\in\operatorname{GL}(2;\mathbb{C}),\quad c\in\mathbb{C}^{3}\right\}$
where $\Phi_{B,d}([x:y:z]_{\omega})=[b_{11}x+b_{12}y:b_{21}x+b_{22}y:z+c_{xx}x^{2}+c_{xy}xy+c_{yy}y^{2}]_{\omega}$ for
$B=\begin{pmatrix}b_{11}&b_{12}\\ b_{21}&b_{22}\end{pmatrix},\quad c=(c_{xx},c_{xy},c_{yy}).$
Note that $\Phi_{B,c}=\Phi_{-B,c}$. Using this group we can assume that the two cusps are at $[1:0:0]_{\omega}$ and $[0:1:0]_{\omega}$ and $z=0$ is the $2$-curve tangent to the cusps and one of $\mathbb{E}_{6}$ points is $[1:1:1]_{\omega}$. The coordinates of the other $\mathbb{E}_{6}$ need to be computed. Note that $[x:y:z]_{\omega}\mapsto[y:x:z]_{\omega}$ is the only automorphism fixing this family of curves.
Remark 6.2.
Altough the above approach is quite natural, computations become too heavy and they do not end with a solution.
There is an automorphism of $\mathbb{P}^{2}_{\omega}$ sending the above family of curves to curves satisfying:
  • the $\mathbb{E}_{6}$ points are at $[1:0:0]_{\omega}$ and $[0:1:0]_{\omega}$;
  • the cusps are at $[1:1:0]_{\omega}$ and $[a_{1}:1:1]_{\omega}$;
  • the $2$-curve tangent to the cusps is $z=bxy$ for some $b$;
  • as with the previous family the are fixed by $[x:y:z]_{\omega}\mapsto[y:x:z]_{\omega}$.
The conditions about the singular points give a system of equations. Direct attempts failed and in the notebook OcticInvolution of Sagemath we obtain the existence of a unique solution up to automorphism. We have normalized this solution to have a simpler form.
Theorem 6.3.
Let $C_{8,2}$ be a projective plane curve of degree $8$ having $6$ singular points of type $\mathbb{E}_{6}$ and fixed by an involution. Then it is projectively equivalent to the curve of equation
$-\frac{11}{3}x^{5}y^{3}-\frac{407}{16}x^{4}y^{4}-44x^{3}y^{5}-\frac{11}{8}x^{4}y^{2}z^{2}+\frac{33}{2}x^{2}y^{4}z^{2}+\frac{27}{176}x^{4}z^{4}$
$-\frac{4}{11}x^{3}yz^{4}-\frac{49}{11}x^{2}y^{2}z^{4}-\frac{48}{11}xy^{3}z^{4}+\frac{243}{11}y^{4}z^{4}-\frac{5}{6}x^{2}z^{6}+10y^{2}z^{6}+z^{8}$ $=0$
This curve is not fixed by any other automorphism.
The proof of the unicity relies on the Sagemath worksheet, but the fact that this equation satisfies the condition is much easier, see CheckCurveInvolution.
Theorem 6.4.
The fundamental group of the complement of $C_{8,2}$ is
$G_{2}=\langle x,y,z\mid[x,z]=1,\ xyx=yxy,\ yzy=zyz,\ (xy^{2}z)^{2}=1\rangle,\quad$
$G_{2}/G_{2}^{\prime}\cong\mathbb{Z}/8,\quad G_{2}^{\prime}/G_{2}^{\prime\prime}\cong\mathbb{Z}/3,\quad G_{2}^{\prime\prime}/G_{2}^{\prime\prime\prime}\cong(\mathbb{Z}/2)^{4}\quad G_{2}^{\prime\prime\prime}\cong\mathbb{Z}^{3}\times\mathbb{Z}/2.$
In particular it is not isomorphic to the fundamental group in Proposition 4.1, and hence $C_{8,2}$ is not isotopic to the symplectic curve in §4.
This theorem has been proved using Sagemath and Sirocco, see the details in the notebook FundamentalGroupInvolution. Note that Sirocco uses interval arithmetic which certifies the results.
7. Algebraic curves with $\mathbb{Z}/3$-action
From the lemmas of §5 we may assume that the automorphism of order $3$ is $\Phi_{3}:\mathbb{P}^{2}\to\mathbb{P}^{2}$, $\Phi_{3}([x:y;z])=[\zeta x:\overline{\zeta}y:z]$ where $\zeta$ is a primitive cubic root of unity. Let $C_{8,3}\in\Lambda_{\text{\rm alg}}$ fixed by the $\Phi_{3}$. Let $X_{3}:=\mathbb{P}^{2}/\Phi_{3}$ its quotient and let $D_{8,3}\subset X_{3}$ be the image of $C_{8,3}$. The surface $X_{3}$ is normal with three isolated cyclic points of type $\frac{1}{3}(1,-1)$. This notation stands for the following. Let $\mu_{d}$ be the group of $d$-roots of unity in $\mathbb{C}$. Then $\frac{1}{d}(a,b)$ is the quotient of $\mathbb{C}^{2}$ by the action of $\mu_{d}$ defined by $\zeta\cdot(x,y)=(\zeta^{a}x,\zeta^{b}y)$.
[…]
Figure 5. Surface $X_{3}$ with the image of the curve and $(1,1)-$blow-up of (B1).
There is a birational transformation to pass from $X_{3}$ to $\mathbb{P}^{2}$. These are the steps:
  1. (B1)
    $(1,1)$-blow-up of the image of $[0:0:1]$ in $X_{3}$, with exceptional component $E$. We obtain a singular ruled surface with four singular points in two fibers, the strict transforms of $L_{x},L_{y}$. The new ones are of type $\frac{1}{3}(1,1)$. The two sections in the right-hand side of Figure 5 have self-interesection $\frac{1}{3}$ ($L_{z}$ below) and $-\frac{1}{3}$ ($E$ above), see [AMO14] for details on weighted blow-ups.
  2. (B2)
    $(1,1)$-blow-up of the two points of type $\frac{1}{3}(1,1)$, with exceptional components $E_{x},E_{y}$ of self-interesection $-3$. The self-interesection of the strict transforms of $L_{x},L_{y}$ is $-\frac{1}{3}$.
  3. (B3)
    Blow-down of the strict transforms of the images of the lines $x=0$, $y=0$; it is the inverse of a $(1,3)$-blow-up of a smooth points. The result is a smooth surface, actually the Hirzebruch ruled surface $\Sigma_{1}$ where the $(-1)$-curve is $E$.
  4. (B4)
    Contract the $(-1)$-curve $E$.
[…]
Figure 6. $(1,1)$-Blow-ups of (B2), blow-downs of (B3) and blow-down of (B4).
Actually all this operation has simple coordinates. The composition of the quotient and the birational map is a rational map $\Theta:\mathbb{P}^{2}\dashrightarrow\mathbb{P}^{2}$ given by
$\Theta([x:y:z])=[x^{3}:y^{3}:xyz].$
Lemma 7.1.
Let $C$ the image of $C_{8,3}$ by $\Theta$. Then $C$ is a curve of degree $8$, with two singular points $\mathbb{E}_{6}$ and two singularities with the topological type of $(u-v^{2})(u^{2}-v^{6})=0$ having maximal contact with the tangent line.
We proceed as in §6. Let us take
$f(x,y,z)=\sum_{i+j+k=8}a_{ijk}x^{i}y^{j}z^{k}.$
with $a_{008}=1$ since $[0:0:1]\notin C$. We place the reducible singular points at $[1:0:0]$ and $[0:1:0]$ with respective tangent lines $y=0$ and $z=0$. One of the $\mathbb{E}_{6}$ points is at $[1:1:1]$ and for the other one we use two new variables. The only automorphism fixing this family of curves is $[x:y:z]\mapsto[y:x:z]$.
The system of equations is more complicated that the one in §6, but we managed to obtain the solutions using Sagemath, see the notebook OcticAuto3. In Appendix A, the common procedure is explained. To describe the solution we need to introduce the number field $\mathbb{K}:=\mathbb{Q}[\eta]$, where $\eta$ is a solution of $p(t):=t^{4}-2t^{3}+t^{2}-2t-2$. This polynomial has two real roots $\eta_{1},\eta_{2}$ and two complex conjugate roots $\eta_{3},\eta_{4}$.
Theorem 7.2.
Let $C_{8,3}$ be a projective plane curve of degree $8$ having $6$ singular points of type $\mathbb{E}_{6}$ and fixed by an automorphism of order $3$. Then it is projectively equivalent to a curve $C^{\eta_{i}}_{8,3}$ whose equation is obtained as follows. Let
$G_{0}(x,y,z)=\frac{F(x^{3},y^{3},xyz)}{x^{8}y^{8}},$
where $F$ is the equation in the Appendix B. Then, $G(x,y,z):=G_{0}(x+\zeta y,x+\overline{\zeta}y,z)$ with coefficients in $\mathbb{K}=\mathbb{Q}[\eta_{i}]$. This curve is not fixed by any other automorphism.
The fundamental group of the complement of any such curve is cyclic of order $8$.
The proof of this theorem can be checked in OcticAuto3. The computation of the fundamental group takes much longer than it took in the case of §6 and it has been done with Sagemath and Sirocco, see the notebook FundamentalGroupAuto3.
As for the other type of curves, the long computation is only needed to prove that these curves are the only ones. It is easier to prove that the satisfy the required condition, see CheckCurveAuto3.
8. Alternative way to compute the fundamental groups
There is an alternative way to compute this fundamental group. We can compute $G_{3}^{\textrm{orb}}:=\pi_{1}^{\textrm{orb}}(X_{3}\setminus D_{8,3}^{\eta_{i}})$ and $G_{2}^{\textrm{orb}}:=\pi_{1}^{\textrm{orb}}(X_{2}\setminus\tilde{C}_{8,2})$. In this particular situation it does not really save computation time but in other cases it allows to obtain a faster and computer-free approach.
The orbifold fundamental group $\pi_{1}^{\textrm{orb}}(X_{2}\setminus\tilde{C}_{8,2})$ is computed following several steps, see Alternatives2:
  1. (Orb21)
    Blow up $[0:0:1]_{\omega}$; we obtain a surface $\Sigma_{2}$ (a ruled Hirzebruch surface) with an exceptional component $E$, with self-interesection $-2$. We compute the group $\pi_{1}(\Sigma_{2}\setminus(\tilde{C}_{8,2}\cup L_{z}\cup E))$.
  2. (Orb22)
    To compute this group we consider an affine chart, say the complement of $E$ and $L_{x}$, using the standard Zariski-van Kampen method. In Alternatives2 we have a finitely presented group with five generators $x_{0},\dots,x_{4}$, where $x_{2}$ is a meridian of $L_{z}$ and $e:=(x_{0}\cdot\ldots\cdot x_{4})^{-1}$ is a meridian of $E$. Following [KK03], a meridian of $L_{x}$ is $e^{2}$.
  3. (Orb23)
    The group $G^{2}_{\textrm{orb}}$ is obtained by adding the relations $x_{2}^{2}=e^{2}=1$.
  4. (Orb24)
    The group $G_{2}$ is the kernel of the map $G_{2}^{\textrm{orb}}\twoheadrightarrow\mathbb{Z}/2$ defined by $x_{i}\mapsto 0$, $i\neq 2$, and $x_{2}\mapsto 1$. In Alternatives2, we prove that $x_{2}$ is central and of order $2$. Hence $G_{2}^{\textrm{orb}}\cong G_{2}\times\mathbb{Z}/2$.
  5. (Orb25)
    Actually $G_{2}$ is the orbifold fundamental group of the complement of $\tilde{C}_{8,2}$, where the unique orbifold point is the singular one.
We follow a similar strategy to compute the orbifold fundamental groups $G_{3}^{\textrm{orb}}=\pi_{1}^{\textrm{orb}}(X_{3}\setminus\tilde{D}_{8,3}^{\eta_{i}})$, see Alternatives3:
  1. (Orb31)
    We start with the final birational model of the rational map and compute $\pi_{1}(\mathbb{P}^{2}\setminus(\tilde{D}_{8,3}^{\eta_{i}}\cup E_{x}\cup E_{y}))$. Actually, we take the affine chart of the complement of $E_{x}$ and compute the fundamental group of the complement of $\tilde{D}_{8,3}^{\eta_{i}}$ and $E_{y}$.
  2. (Orb32)
    Using the standard Zariski-van Kampen method we obtain in Alternatives3 a finitely presented group with nine generators $x_{0},\dots,x_{8},\ell_{y}$, where $e_{y}$ is a meridian of $E_{y}$, the $x_{i}$’s are meridians of $\tilde{D}_{8,3}^{\eta_{i}}$ $e:=(x_{0}\cdot\ldots\cdot x_{8})^{-1}$ is a meridian of $E$, and $e_{x}:=e_{y}^{-1}\cdot e$ is a meridian of $E_{x}$.
  3. (Orb33)
    Following [Mum61], we deduce that for the group $G_{3}^{\textrm{orb}}$ we have to add the relations deduced from the divisor $E+E_{x}+E_{y}$ in Figure 7:
    $e_{x}\cdot e_{y}=e\ (\text{known}),\quad e=e_{x}^{3}=e_{y}^{3}\Rightarrow e=e_{x}\cdot e_{y}=e_{x}^{3}=1.$
  4. (Orb34)
    With this new relation we have computed in Alternatives3 that all the groups are $\mathbb{Z}/24$ and hence we recover the abelianity of $G_{3}$.
[…]
Figure 7. Minimal resolution of $\hat{X}_{3}$
In Alternatives3 we have computed a simplified braid monodromy for $\tilde{D}_{8,3}^{\eta_{i}}$ which may give some hints about the topological equivalence of these curves.
9. Conclusions
We summarize the results an open questions.
  1. (C1)
    There is no homeomorphism $\Phi_{i}:\mathbb{P}^{2}\to\mathbb{P}^{2}$ such that $\Phi_{i}(C_{8,3}^{\eta_{i}})=C_{8,2}$.
  2. (C2)
    There is no homeomorphism $\Psi_{i}:\mathbb{P}^{2}\to\mathbb{P}^{2}$ such that $\Psi_{i}(C_{8,3}^{\eta_{i}})=C_{\text{\rm symp}}$.
  3. (C3)
    There is no homeomorphism $\Psi_{i}:\mathbb{P}^{2}\to\mathbb{P}^{2}$ such that $\Psi_{i}(C_{8,2})=C_{\text{\rm symp}}$.
  4. (C4)
    The complex conjugation is a homeomorphism $\Phi:\mathbb{P}^{2}\to\mathbb{P}^{2}$ such that $\Phi(C_{8,3}^{\eta_{3}})=C_{8,3}^{\eta_{4}}$.
  5. (C5)
    The existence of homeomorphisms $\Phi_{i,j}:\mathbb{P}^{2}\to\mathbb{P}^{2}$ such that $\Phi(C_{8,3}^{\eta_{i}})=C_{8,3}^{\eta_{j}}$ is an open question, for $i\neq j$ and $\{i,j\}\neq\{3,4\}$.
  6. (C6)
    The existence of a homeomorphism $\Phi_{3,4}:\mathbb{P}^{2}\to\mathbb{P}^{2}$ such that $\Phi(C_{8,3}^{\eta_{3}})=C_{8,3}^{\eta_{4}}$ which preserves the orientation of the curves is an open question.
  7. (C7)
    The existence of other curves in $\Lambda_{\text{\rm alg}}$ is an open question.
  8. (C8)
    The existence of curves in $\Lambda_{\text{\rm alg}}$ isotopic to $C_{\text{\rm symp}}$ is an open question.
10. Perspectives
A direct approach to compute $\Lambda_{\text{\rm alg}}$ seems to be hopeless. Isolating special properties for the known solutions would help to get new ideas that would allow either to discard new cases or to obtain some new ones. We know that there is no more curve in $\Lambda_{\text{\rm alg}}$ fixed by a non-trivial homeomorphism.
In particular we are going to compute the smooth quartics of §5. Let us consider the the $2$-dimensional projective system formed by the closure of the family of quintics having ordinary double points at the six points $P_{1},\dots,P_{6}$. The intersection number of two such quintics at the base points is at least $24$, so, they intersect at another point. If we blow up the six points the strict transforms of the quintics are smooth rational curves with self-intersection $-1$.
In this closure we find also the curves formed by $\mathcal{C}_{i}$ and a line passing through $P_{i}$, which are exceptional elements of the family. Their strict transforms are disjoint from the strict transforms of the irreducible quintics in the system. The Cremona transformation described in Proposition 5.7 is obtained by blowing-down the strict transforms of these conics.
The notebooks Birational2 and Birational3 contain the computations leading to the following results. Moreover the systems of points described in Proposition 5.7 are also computed.
Theorem 10.1.
The curve $C_{8,2}$ is birationally equivalent to
$z^{4}-3x^{2}z^{2}+y^{2}z^{2}-36x^{3}y+45x^{2}y^{2}-12xy^{3}=0.$
Theorem 10.2.
The curve $C_{8,3}^{\eta_{i}}$ is birationally equivalent to
$z^{4}+\frac{3}{38}b_{12}xyz^{2}+\frac{1}{19}(2b_{01}+\zeta c_{01})x^{3}z+\frac{1}{19}(2b_{01}+\overline{\zeta}c_{01})y^{3}z+\frac{3}{19}b_{20}x^{2}y^{2}=0,$
where
$b_{12}$ $=-97\eta_{i}^{3}-23\eta_{i}^{2}-130\eta_{i}-92$
$b_{01}$ $=74\eta_{i}^{3}+6\eta_{i}^{2}+109\eta_{i}+75$
$c_{01}$ $=-51\eta_{i}^{3}+\eta_{i}^{2}-42\eta_{i}-35$
$b_{20}$ $=3596\eta_{i}^{3}+585\eta_{i}^{2}+4862\eta_{i}+3325.$
Appendix A Strategy of the computations
In §6 and §7 we need to find the zero locus of an ideal $J_{0}$ in a ring $\mathbb{C}[a_{1},\dots,a_{n}]$. More precisely we look for non-degenerate solutions, since the conditions imposed are closed conditions and the space we are looking for is only locally-closed.
The existence of degenerate solutions is a big computational problem. The strategy followed consists to define a tree of ideals whose root is $J_{0}$. This tree has levels and at each level we eliminate a variable.
Let us assume that we have inductively constructed un ideal $J_{j,k}\subset\mathbb{C}[a_{1},\dots,a_{n-k}]$. Using heuristic arguments we choose a generator $f_{0}$ of the ideal and a variable, say $a_{n-k}$, and we compute the resultants with respect to $a_{n-k}$ of $f_{0}$ with the other generators. We factorize each one of these resultants and we eliminate the factors which are known to provide degenerate solutions. With the remaining factors, we combine them to give a family of ideals $J_{j^{\prime},k+1}$ in $\mathbb{C}[a_{1},\dots,a_{n-k}]$.
Some of the leaves of this tree will stop with no solution and we pay attention to the ones ending in prime ideals of $\mathbb{C}[a_{1}]$.
Fix one of these leaves. Actually these ideals have coefficients in $\mathbb{Q}$. A prime ideal $J_{i^{\prime},n-1}\subset\mathbb{C}[a_{1}]$ determines an extension $\mathbb{L}_{1}$ of $\mathbb{Q}$ where a solution leaves. Replacing the value of $a_{1}$ by this solution in the ideal $J_{i,n-2}$ we obtain a new ideal in $\mathbb{L}_{1}[a_{2}]$. We factorize these principal ideals. Either some of these process stop with no solution or we end with one solution.
In both cases we end with only one algebraic solution. For the case of §6 the solution lives in a degree $2$ extension of $\mathbb{Q}$ but the symmetry allows us to end with a rational solution. In the case of §7 the solution lives in $\mathbb{K}_{1}=\mathbb{Q}[\eta,\zeta]$, extension of $\mathbb{Q}$ of degree $8$. The symmetry allows us to end with a solution $\mathbb{K}_{1}=\mathbb{Q}[\eta]$, extension of $\mathbb{Q}$ of degree $4$.
Appendix B Equations
Let $\mathbb{K}_{1}:=\mathbb{Q}[\eta,\zeta]$ and let $\sigma$ be the non-trivial automorphism of $\mathbb{K}_{1}$; the field $\mathbb{Q}[\eta]$ is the fixed field by $\sigma$. When $\eta$ is real, $\sigma$ is the complex conjugation. A curve of Lemma 7.1 is of the form
$F(x,y,z)=F_{0}(xy,z)+2xyz(xF_{1}(xy,z)+yF_{1}^{\sigma}(xy,z))+x^{2}y^{2}(x^{2}F_{2}(xy,z)+y^{2}F_{2}^{\sigma}(xy,z)),$
where $F_{0},F_{1},F_{2}\in\mathbb{K}[t,z]$. We have
$F_{0}(t,z)=$ $z^{8}+\frac{2r_{16}}{19}tz^{6}++\frac{3r_{24}}{19^{2}}t^{2}z^{4}+\frac{2r_{32}}{19^{2}}t^{3}z^{2}+\frac{4r_{40}}{19}t^{4},$
$r_{16}$ $=437\eta^{3}-1270\eta^{2}+1130\eta-1696$
$r_{24}$ $=-596956\eta^{3}+1619007\eta^{2}-1523682\eta+2184414$
$s_{23}$ $=-2064411\eta^{3}+5739587\eta^{2}-5326476\eta+7777170$
$s_{40}$ $=11524593\eta^{3}-28834395\eta^{2}+28396048\eta-38303610.$
$F_{1}(t,z)=$ $\frac{r_{15}+\zeta s_{15}}{19^{3}}z^{4}+\frac{r_{23}+\zeta s_{23}}{19^{2}}tz^{2}+6\frac{r_{31}+4\zeta s_{31}}{19^{2}}t^{2},$
$r_{15}$ $=-157924\eta^{3}+308331\eta^{2}-356378\eta+387894$
$s_{15}$ $=182695\eta^{3}-547611\eta^{2}+485700\eta-752178$
$r_{23}$ $=1276065\eta^{3}-3104444\eta^{2}+3107094\eta-4100620$
$s_{23}$ $=-2064411\eta^{3}+5739587\eta^{2}-5326476\eta+7777170$
$r_{31}$ $=-5295773\eta^{3}+14400235\eta^{2}-13528408\eta+19435018$
$s_{31}$ $=6353433\eta^{3}-16472958\eta^{2}+15895154\eta-22035984$
$F_{2}(t,z)=$ $\frac{r_{22}+2\zeta s_{22}}{19}z^{2}+2\frac{r_{30}+4\zeta s_{30}}{19}t.$
$r_{22}$ $=74354\eta^{3}-196839\eta^{2}+187718\eta-264358$
$s_{22}$ $=138989\eta^{3}-356263\eta^{2}+346016\eta-475522$
$r_{30}$ $=-8288405\eta^{3}+21480135\eta^{2}-20732048\eta+28731618$
$s_{30}$ $=-2845567\eta^{3}+7360179\eta^{2}-7111716\eta+9841218$